OpenStudy (anonymous):

Write this equation in standard form and find the zeros: f(x) = (X²-4)(x²-9)(x-3)

OpenStudy (anonymous):

\[(x^2-4) = (x+2)(x-2)\]\[(x^2-9)=(x+3)(x-3)\] the "zeros" are the inverse of what yu have: -2,2,-3,3

OpenStudy (anonymous):

no. i dont think you did it right